A volatile compound with a mass of 0.8822 grams is placed in an evacuated 0.250 L flask. The flask is heated to evaporate the compound. At 99C, the pressure in the flask is 1.25 atm. What is the molar mass of the compound?


1
Expert's answer
2021-05-11T05:51:20-0400

Using the gas equation,


"PV= nRT"


"1.25 \times 0.250 = \dfrac{0.8822}{x} \times 0.083 \times372"

"0.3125 = \dfrac{0.8822}{x} \times 30.876"


"x = \dfrac{27.23}{0.3125}"


"x = 87.16"